OpenStudy (anonymous):

Luke has a compost pit in his garden. The pit is 6 feet long, 3 feet wide, and 4 feet deep. She plans to make it 1.5 feet deeper. What will be the increase in the volume of the pit? Answer 27 cubic feet 36 cubic feet 99 cubic feet 108 cubic feet

OpenStudy (anonymous):

given; original: l=6 ;w=3; h=4 plan of increase; l=6; w=3 ; h=4+1.5=5.5 solution= (volume of the plan of increase - volume of the original) V =(6)(3)(5.5) - (6)(3)(4) = ? solve. . .

OpenStudy (anonymous):

27 cubic feet